Carrier Management assignment. Fill in the incomplete amounts
Fleet Activity Summary Loaded Miles Empty miles Toal Miles (utilization) Gross linehaul revenue Loaded rate per mile Net revenue per mile 2200 300 $1.70 Estimated Profit and Loss Linehaul Revenue Per mile Tractor Total Variable Costs Driver Wages per mile Fringe benefits (%) Fuel cost per mile $0.48 25% $0.42 Maintenance cost per mile $0.11 $0.10 Insurance cost per mile Total Variable Cost Fixed Cost Tractor Depreciation Trailer depreciation Legalization Corporate overhead Total Fixed Cost Grand Total Cost Per Week $260.00 $42.00 $192.31 $117.02 Operating Profit Profit Margin
Answer #1

Total miles= 2200+300=2500

Gross linehaul revenue= 2500*1.70= $4250

Net revenue per mile= $4250-(300*1.70)=$3740

ESTIMATED PROFIT& LOSS

Particulars amt($) amt($)
per mile Tractor total
Linehaul revenue 1.70 3740
Variable cost:
Driver's wage 0.48 1200
Fringe benefits 300
Fuel cost per mile 0.42 1050
maintenance cost per mile 0.11 275
insurance cost per mile 0.10 250
TOTAL VARIABLE COST 3075
Fixed Cost:
Tractor Depreciation 260
Trailer Depreciation 42
Legalization 192.31
Corporate overhead 117.02
TOTAL FIXED COST 611.33
GRAND TOTAL COST 3686.33
OPERATING PROFIT 53.67
PROFIT MARGIN 1.43%